OpenStudy (anonymous):

A gallon of paint costs $12.99 and covers 400 square feet. How many gallons are needed to paint two coats on the walls and ceiling (not the floor) of a rectangular room that is 30 feet long, 18 feet wide, and 8 feet high? Round to the next whole gallon.

OpenStudy (anonymous):

total areas is \[2\times 18\times 8+2\times 30\times 8=768\] and you have to cover this twice so you need to cover 1536 square feet. divide by 400, round up

OpenStudy (anonymous):

btw it does not ask how much this cost, so the price per gallon is not important here